# Can stellar physics explain GW231123?

### v1

@@ Abstract
The gravitational wave event GW231123 detected by the LIGO interferometers during their fourth observing run features two black holes with source-frame masses of $137^{+22}_{-17}{\rm M}_{\odot}$ and $103^{+20}_{-52}{\rm M}_{\odot}$ — well within or above the pair-instability black hole mass gap predicted by standard stellar evolution theory. Both black holes are also inferred to be rapidly spinning ($\chi_{1}\simeq 0.9$, $\chi_{2}\simeq 0.8$). The primary object in GW231123 is the heaviest stellar mass black hole detected to date, which, together with its extreme rotation, raises questions about its astrophysical origin. Accounting for the unusually large spin of $\sim 0.9$ with hierarchical mergers requires some degree of fine tuning. We investigate whether such a massive, highly spinning object could plausibly form from the collapse of a single rotating massive star. We simulate stars with an initial core mass of $160\,{\rm M}_{\odot}$ — sufficient to produce BH masses at the upper edge of the 90% credible interval for $m_{1}$ in GW231123 — across a range of rotation rates and $^{12}\mathrm{C}(\alpha,\gamma)^{16}\mathrm{O}$ reaction rates. We find that: (i) rotation shifts the pair-instability mass gap to higher masses, introducing a significant ingredient that correlates masses and spins in gravitational wave predictions; and (ii) highly spinning BHs with masses $\gtrsim 150\rm M_{\odot}$ can form above the mass gap, implying that stellar evolution alone is sufficient to explain GW231123. Our results suggest that the primary object of GW231123 may be the first directly observed black hole that formed via direct core collapse following the photodisintegration instability.
preprint: IPPP/25/52
Introduction
@@
$m_{1}=137^{+22}_{-17}\,{\rm M}_{\odot},\quad m_{2}=103^{+20}_{-52}\,{\rm M}_{\odot},$ (1)
@@
Stellar evolution theory predicts a “pair‐instability” black hole mass gap (BHMG) between approximately $60$ and $130\,{\rm M}_{\odot}$, driven by electron–positron pair production in helium cores that powers (pulsational) pair‐instability supernovae (hereafter (P)PISN) and completely disrupts stars with core masses up to $\sim 135{\rm M}_{\odot}$ [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20]. Various stellar parameters and other effects determine the location of the BHMG, most importantly the $^{12}\mathrm{C}(\alpha,\gamma)^{16}\mathrm{O}$ rate [15, 21, 16, 17, 18], which can give variations of $\sim 47{\rm M}_{\odot}$ [17] in the location of the upper and lower edges of the BHMG.
@@
If it had zero spin, the primary BH in GW231123 would likely lie above the BHMG unless the $^{12}\mathrm{C}(\alpha,\gamma)^{16}\mathrm{O}$ reaction rate were $\sim 3\sigma$ smaller than its median value [17]. However, this is based on simulations of (P)PISN which have neglected rotation, so this conclusion is unlikely to apply to GW231123 and its extremely high spins. Indeed, previous theoretical work [22] has shown that rotation can raise the lower edge of the gap by $\sim 15\%$ (from $45.5{\rm M}_{\odot}$ to $52.4{\rm M}_{\odot}$). It is currently not known how rotation affects the upper edge of the BHMG. These considerations raise the question of whether stellar evolution can produce a BH as massive and highly spinning as the primary in GW231123.
Core collapse instabilities
@@
The predicted BHMG is a direct result of the physics of pair‐instability in massive helium cores. At core temperatures $\gtrsim 10^{9}\,\mathrm{K}$, high‐energy photons convert into non-relativistic electron–positron pairs, stealing pressure support and causing the core to contract. In turn, this contraction ignites explosive nuclear burning, primarily of oxygen. If the released nuclear energy exceeds the gravitational binding energy, the star is completely disrupted with no BH remnant. In a slightly lower mass range the explosion is weaker and instead ejects the outer layers in pulses (a PPISN), removing mass until the remaining core collapses to a lighter BH. Above a critical helium core mass ($\sim 130{\rm M}_{\odot}$), stellar cores become so hot ($T_{c}\gtrsim 9\times 10^{9}\,\mathrm{K}$) that high‐energy photons begin to photodisintegrate iron‐group nuclei into $\alpha$‐particles and free nucleons. This endothermic process robs the core of pressure support, accelerating collapse — a phenomenon known as the photodisintegration instability. In this scenario, the star implodes directly to a BH, defining the upper edge of the BHMG.
As noted above, the effects of rotation on the location of the upper edge of the BHMG are unknown. In general, the extra support provided by the centrifugal force results in lower central densities and temperatures compared with non-rotating stars [22, 23], so it can be expected that rotation delays the onset of photodisintegration instability and shifts the upper edge of the BHMG to heavier masses.
@@
As the $^{12}\mathrm{C}(\alpha,\gamma)^{16}\mathrm{O}$ rate is raised, a larger fraction of 12C is converted into 16O during core helium burning. This higher oxygen abundance results in a stronger explosion, whilst the lower 12C abundance implies less carbon is available to form a convective 12C burning shell to counteract contraction [16]. Consequently, both the lower and upper edges of the mass gap move downward in tandem, preserving a nearly constant gap width of $\Delta M_{\rm BHMG}\approx 80^{+9}_{-5}{\rm M}_{\odot}$ [17]. Across the full $\pm 3\sigma$ uncertainty in the reaction rate, the lower boundary shifts from $\sim 59^{+34}_{-13}{\rm M}_{\odot}$ down to lighter values, while the upper boundary moves from $\sim 139^{+30}_{-14}{\rm M}_{\odot}$ to correspondingly lower masses [17].
Simulation suite
To assess the impact of rotation and nuclear physics on the boundaries of the pair-instability regime, we compute a grid of stellar models with initial helium core mass $160{\rm M}_{\odot}$. This value was chosen such that, if no PISN occurs, the star can be expected to form a BH with mass at the edge of the measured $90\%$ confidence for $m_{1}$ in GW231123. We vary the stellar rotation rate $\Omega/\Omega_{\mathrm{crit}}$ between 0 and $1$ to capture the effects of increasing centrifugal support. The critical angular velocity
@@
is defined as the rotation rate at which the outward centrifugal acceleration at the stellar equator exactly balances the inward effective gravitational acceleration, where $M$ is the gravitational mass, $R_{\rm eq}$ the equatorial radius, and $\Gamma\equiv L/L_{\rm Edd}=L\kappa/{4\pi GMc}$ is the local Eddington factor which accounts for radiation pressure. Additionally, we explore the variation of the rate of the temperature-dependent $^{12}\mathrm{C}(\alpha,\gamma)^{16}\mathrm{O}$ reaction over $\pm 3\sigma$ from its median rate $R_{\rm med}(T)$. Here, $\sigma$ parametrizes the variation as $R_{\sigma}(T)=R_{\rm med}(T)\exp[\sigma\mu(T)]$ with $\mu(T)$ the uncertainty at each temperature point, which is assumed to follow a log-normal distribution [24, 16, 17].
Our simulations were performed using the stellar structure code MESA, version 15140 [29, 30, 31, 32, 33, 34]. While MESA is a one-dimensional code — meaning that it assumes spherical symmetry — it is capable of simulating rotating objects via the shellular approximation, where the radial coordinate $r$ is replaced with the volume-equivalent radius of an isobar. A comprehensive description of rotation in MESA is provided in Ref. [33]. The technical details of our simulations are given in the End Matter, and our code will be made public upon publication of this work. We do not include the Spruit-Tayler (ST) dynamo [35], which allows for efficient differential rotation and retention of angular momentum throughout the evolution. The ST dynamo is commonly invoked in stellar models to enforce near-rigid rotation by coupling angular momentum between stellar layers [36, 37]. However, its physical validity remains debated, with some studies raising doubts about both the efficiency and even the onset of the magnetic instabilities required to close the dynamo loop under realistic stellar conditions (e.g., [38, 39]). The high spins inferred for both components of GW231123 are difficult to reconcile with models that include the ST dynamo, which predict nearly non-rotating BHs [22]; this suggests that angular momentum transport in massive stars may be much less efficient than the ST prescription assumes.
@@
Figure 1: Final BH mass as a function of initial rotation with initial helium core mass $160\rm{\rm M}_{\odot}$, for several values of the $^{12}\mathrm{C}(\alpha,\gamma)^{16}\mathrm{O}$ rate. The top panel shows results as a function of the initial stellar rotation $\Omega$ in units of the critical value $\Omega_{\rm crit}$; the bottom panel shows results as a function of the BH spin $\chi$. Stars with faster rotation and smaller $^{12}\mathrm{C}(\alpha,\gamma)^{16}\mathrm{O}$ reaction rate undergo PISN and are not shown in the figure. The shaded contours indicate the $90\%$ credible intervals of the primary mass (orange) and spin (purple) in GW231123 [1]
@@
Some simulations feature spin parameters exceeding unity at core-collapse. Although the BH may form with maximal spin, further accretion can reduce its spin if the added mass is not accompanied by sufficient angular momentum, or if excess angular momentum is expelled by feedback to avoid violating the Kerr limit. To estimate the final spin of the BH, we adopt a version of the model developed in Ref. [40], consistent with previous work [22]. We assume that the innermost $3{\rm M}_{\odot}$ of the progenitor collapses promptly into a BH with maximal spin. We have explicitly verified that the results are insensitive to the size of this core, see End Matter for more details. The remainder of the star is assumed to either fall directly into the BH or accrete through a disk, delivering additional mass and angular momentum. Accretion is regulated by feedback, ensuring that the final spin does not exceed the Kerr limit ($\chi\leq 1$).
Impact of rotation on the black hole mass gap

### v2

# Can GW231123 have a stellar origin?

@@ Abstract
The gravitational wave event GW231123 detected by the LIGO interferometers during their fourth observing run features two black holes with source-frame masses of $137^{+23}_{-18}{\rm M}_{\odot}$ and $101^{+22}_{-50}{\rm M}_{\odot}$in the range of the pair-instability black hole mass gap predicted by standard stellar evolution theory. Both black holes are also inferred to be rapidly spinning ($\chi_{1}\simeq 0.9$, $\chi_{2}\simeq 0.8$). The primary object in GW231123 is the heaviest stellar mass black hole detected to date, which, together with its extreme rotation, raises questions about its astrophysical origin. Accounting for the unusually large spin of $\sim 0.9$ with hierarchical mergers requires some degree of fine tuning. We investigate whether such a massive, highly spinning object could plausibly form from the collapse of a single rotating massive star. We simulate stars with an initial core mass of $160\,{\rm M}_{\odot}$ — sufficient to produce BH masses at the upper edge of the 90% credible interval for $m_{1}$ in GW231123 — across a range of rotation rates and $^{12}\mathrm{C}(\alpha,\gamma)^{16}\mathrm{O}$ reaction rates.  We allow for differential rotation to explore the high-spin regime.  In this limit of weak angular momentum transport, we find that: (i) rotation shifts the pair-instability mass gap to higher masses, introducing an important correlation between masses and spins in gravitational wave predictions; and (ii) highly spinning BHs with masses $\gtrsim 150\rm M_{\odot}$ can form above the mass gap. Our results suggest that the primary object of GW231123 may be the first directly observed black hole that formed via direct core collapse following the photodisintegration instability.
keywords:
black hole mergers — gravitational waves
pubyear: 2026pagerange: Can GW231123 have a stellar origin?LABEL:lastpage
1 Introduction
@@
$\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}m_{1}=137^{+23}_{-18}\,{\rm M}_{\odot},\quad m_{2}=101^{+22}_{-50}\,{\rm M}_{\odot},$} (1)
@@
Ostensibly, the combination of large masses and large spins in GW231123 could be explained by binaries formed through hierarchical mergers (for context, see Gerosa and Fishbach 2021 and references therein) and this appears to be a leading explanation in some of the recent litterature (Abac and others, 2025; Stegmann et al., 2025; Li et al., 2025; Paiella et al., 2025). We argue that a hierarchical merger origin for GW23112’s primary BH is unlikely; cf. Bartos and Haiman (2026) for a similar argument. Hierarchical mergers in dense stellar environments produce a spin distribution that is sharply peaked at $\chi\sim 0.7$ (Gerosa and Berti, 2017; Fishbach et al., 2017), which matches the primary BH spin in GW231123 only near the lower edge of its 90% credible interval. Spins of $\chi\sim 0.9$ are too large to be explained as merger remnants (Gerosa et al., 2021), requiring substantial fine‐tuning of the progenitor binary (involving either a somewhat extreme mass ratio and/or mostly aligned spins). Motivated by these considerations, in this work we explore the possibility that the primary component of GW231123 could have formed from the direct collapse of a rotating massive star.
@@
Stellar evolution theory predicts a “pair‐instability” black hole mass gap (BHMG) between approximately $60$ and $130\,{\rm M}_{\odot}$, driven by electron-positron pair production in helium cores that powers (pulsational) pair‐-instability supernovae (hereafter (P)PISN) and completely disrupts stars with core masses up to $\sim 135{\rm M}_{\odot}$ (Fowler and Hoyle, 1964; Barkat et al., 1967; Rakavy and Shaviv, 1967; Fraley, 1968; Spera and Mapelli, 2017; Woosley et al., 2002; Woosley, 2017; Farmer et al., 2019, 2020; Mehta et al., 2022; Farag et al., 2022; Woosley and Heger, 2021; Hendriks et al., 2023). Various stellar parameters and other effects determine the location of the BHMG, most importantly the $^{12}\mathrm{C}(\alpha,\gamma)^{16}\mathrm{O}$ rate (Takahashi, 2018; Farmer et al., 2019; Sakstein et al., 2020; Farmer et al., 2020; Mehta et al., 2022; Farag et al., 2022), which can give variations of $\sim 47{\rm M}_{\odot}$ (Mehta et al., 2022) in the location of the upper and lower edges of the BHMG.
@@
If it had zero spin, the primary BH in GW231123 would likely lie above the BHMG unless the $^{12}\mathrm{C}(\alpha,\gamma)^{16}\mathrm{O}$ reaction rate were $\sim 3\sigma$ smaller than its median value (Mehta et al., 2022). However, this is based on simulations of (P)PISN which have neglected rotation, so this conclusion is unlikely to apply to GW231123 and its extremely high spins. Indeed, previous theoretical work (Marchant and Moriya, 2020) has shown that rotation can raise the lower edge of the gap by $\sim 15\%$ (from $45.5{\rm M}_{\odot}$ to $52.4{\rm M}_{\odot}$). It is currently not known how rotation affects the upper edge of the BHMG.
A major uncertainty in modeling the rotational evolution of very massive stars is the efficiency of internal angular momentum transport. Much of the empirical support for strong core-envelope coupling, often modeled via the Spruit-Tayler (ST) dynamo, comes from low- and intermediate-mass stars. Its operation in very massive, radiation-dominated, and rapidly evolving stars, however, remains largely unconstrained observationally. The high spin inferred for the primary of GW231123 would be difficult to reconcile with standard stellar evolution models that enforce near-rigid rotation throughout the star’s lifetime. Motivated by this tension, we explore an optimistic bracketing scenario in which internal angular momentum transport is inefficient, allowing the core to retain substantial rotation up to collapse. This assumption is not meant to be representative of typical massive stars, but rather to test whether the photodisintegration formation channel can remain viable even in the extreme case required by GW231123. Finally, we note that several studies have proposed evolutionary scenarios in which progenitors retain or regain substantial core rotation, including chemically homogeneous evolution (de Mink and Mandel, 2016; Marchant et al., 2016; Popa and de Mink, 2025), early envelope stripping in binaries (Qin et al., 2018), and tidal spin-up (Olejak and Belczynski, 2021). Our results provide a first indication of how pair-instability boundaries might be modified in such scenarios, independent of the detailed mechanism by which the core is spun up.
2 Core collapse instabilities
@@
The predicted BHMG is a direct result of the physics of pair‐instability, a well-studied effect  (Fowler and Hoyle, 1964; Barkat et al., 1967; Rakavy and Shaviv, 1967; Fraley, 1968; Spera and Mapelli, 2017; Woosley et al., 2002; Woosley, 2017; Takahashi, 2018; Farmer et al., 2019, 2020; Mehta et al., 2022; Farag et al., 2022; Woosley and Heger, 2021; Hendriks et al., 2023) in massive post-main-sequence stars. At core temperatures $\gtrsim 10^{9}\,\mathrm{K}$, high‐energy photons convert into non-relativistic electron-positron pairs, softening the equation of state such that the adiabatic index $\Gamma_{1}<4/3$, causing the core to contract, which in turn leads to explosive nuclear burning, primarily of oxygen.  If the released nuclear energy exceeds the gravitational binding energy, the star is completely disrupted with no BH remnant. In a slightly lower mass range the explosion is weaker and instead ejects the outer layers in pulses (a PPISN), removing mass until the remaining core collapses to a lighter BH. Above a critical helium core mass ($\sim 130{\rm M}_{\odot}$), stellar cores become so hot ($T_{c}\gtrsim 9\times 10^{9}\,\mathrm{K}$) that high‐energy photons begin to photodisintegrate iron‐group nuclei into $\alpha$‐particles and free nucleons. This endothermic process decreases pressure support in the core (relative to continued nuclear burning), accelerating collapse — a phenomenon known as the photodisintegration instability. In this scenario, the star implodes directly to a BH, defining the upper edge of the BHMG.
Rotation impacts pair instability through two related mechanisms. First, the extra support provided by the centrifugal force results in lower temperatures at a given central density compared with non-rotating stars (Marchant and Moriya, 2020; Huynh et al., 2025). This shifts the star away from the regime where the effective adiabatic index drops below the critical value of $4/3$ for dynamical instability. Second, if the core contracts while approximately conserving angular momentum, it spins up, further increasing centrifugal support during contraction, effectively lowering the critical value of $\Gamma_{1}$ for stability below $4/3$. Together, these effects act to stabilize rapidly rotating stars against pair instability. It is presently not known what the effect on photodisintegration instability is.
@@
As the $^{12}\mathrm{C}(\alpha,\gamma)^{16}\mathrm{O}$ rate is raised, a larger fraction of 12C is converted into 16O during core helium burning. This higher oxygen abundance results in a stronger explosion, whilst the lower 12C abundance implies less carbon is available to form a convective 12C burning shell to counteract contraction (Farmer et al., 2020). Consequently, both the lower and upper edges of the mass gap move downward in tandem, preserving a nearly constant gap width of $\Delta M_{\rm BHMG}\approx 80^{+9}_{-5}{\rm M}_{\odot}$ (Mehta et al., 2022). Across the full $\pm 3\sigma$ uncertainty in the reaction rate, the lower boundary shifts from $\sim 59^{+34}_{-13}{\rm M}_{\odot}$ down to lighter values, while the upper boundary moves from $\sim 139^{+30}_{-14}{\rm M}_{\odot}$ to correspondingly lower masses (Mehta et al., 2022).
3 Simulation suite
To assess the impact of rotation and nuclear physics on the boundaries of the pair-instability regime, we computed a grid of stellar models with an initial helium core mass of $160{\rm M}_{\odot}$. This value was chosen such that, if no PISN occurs, the star can be expected to form a BH with mass at the edge of the measured $90\%$ confidence for $m_{1}$ in GW231123. We vary the stellar rotation rate $\Omega/\Omega_{\mathrm{crit}}$ between 0 and $1$ to capture the effects of increasing centrifugal support. The critical angular velocity
@@ 3 Simulation suite
is defined as the rotation rate at which the outward centrifugal acceleration at the stellar equator exactly balances the inward effective gravitational acceleration, where $M$ is the gravitational mass, $R_{\rm eq}$ the equatorial radius, and $\Gamma\equiv L/L_{\rm Edd}=L\kappa/{4\pi GMc}$ is the local Eddington factor which accounts for radiation pressure. Additionally, we explore the variation of the rate of the temperature-dependent $^{12}\mathrm{C}(\alpha,\gamma)^{16}\mathrm{O}$ reaction over $\pm 3\sigma$ from its median rate $R_{\rm med}(T)$. Here, $\sigma$ parametrizes the variation as $R_{\sigma}(T)=R_{\rm med}(T)\exp[\sigma\mu(T)]$ with $\mu(T)$ the uncertainty at each temperature point, which is assumed to follow a log-normal distribution (deBoer and others, 2017; Farmer et al., 2020; Mehta et al., 2022).
Our simulations were performed using the stellar structure code MESA, version 15140 (Paxton et al., 2011; Paxton and others, 2013, 2015, 2018, 2019; Jermyn and others, 2023). While MESA is a one-dimensional code — meaning that it assumes spherical symmetry — it is capable of simulating rotating objects via the shellular approximation, where the radial coordinate $r$ is replaced with the volume-equivalent radius of an isobar. A comprehensive description of rotation in MESA is provided in Paxton and others (2019). Our code can be found in the paper’s reproduction package (Sakstein, 2025) at the following URL: https://zenodo.org/records/16898502.
We simulate $160{\rm M}_{\odot}$ rotating helium cores with metallicity $Z=10^{-5}$ from the zero-age horizontal branch (ZAHB) to their end-state — either core collapse or disruption due to PISN — since their hydrogen envelopes are expected to be stripped by strong stellar winds or binary interactions, leaving behind bare helium cores that determine the subsequent evolution (Heger and Woosley, 2002; Eldridge et al., 2008; Belczynski et al., 2016; Woosley, 2017). Rotation is initialized at the ZAHB. Mass loss to Wolf-Rayet winds are described following the prescription of Brott et al. (2011), with a rotational enhancement described in Paxton and others (2013). Stars that approach critical rotation are subject to additional mass losses following Paxton and others (2013) that remove angular momentum to keep them sub-critical. Convective mixing is implemented following the time-dependent formulation of Marchant et al. (2019); Renzo et al. (2020), which captures rapid changes in convective criteria during phases of dynamical instability but reduces to standard mixing length theory on long time-scales. We take the mixing length efficiency to be $\alpha_{\rm MLT}=2.0$. Convection is determined using the Ledoux criterion with semi-convection efficiency $\alpha_{\rm SC}=1.0$. We employ exponential overshooting with $f=0.01$. The switch from convective mixing to overshooting occurs at a distance $f_{p}H_{P}$ into the convective zone with $f_{p}=0.005$ and $H_{p}$ the pressure scale-height. We adopted the AGSS09 metal fractions (Asplund et al., 2009). These choices follow standard practice in recent studies of massive helium star evolution and pair-instability supernovae (Croon et al., 2020; Sakstein et al., 2020; Straight et al., 2020; Croon et al., 2021; Sakstein et al., 2022; Croon and Sakstein, 2024), including the rotation prescription introduced by Marchant and Moriya (2020). We replaced the default MESA $^{12}\mathrm{C}(\alpha,\gamma)^{16}\mathrm{O}$ rate with the state-of-the-art tabulated values from Mehta et al. (2022), which sample the deBoer and others (2017) rates; again consistent with contemporary works (Mehta et al., 2022; Farag et al., 2022; Croon and Sakstein, 2025). Our resolution is higher than most MESA PPISN studies, following ecent works that suggest such high resolutions are needed to fully resolve the core collapse-PPISN transition (Mehta et al., 2022; Farag et al., 2022; Croon and Sakstein, 2025), although we note that a resolution study on the PISN-photodisintegration boundary has yet to be performed. Specifically, we set delta_lgRho_cntr_limit = 0.001d0 and max_dq = 5d-4 to enforce a smaller time-step and higher spatial resolution.
As noted above, our study is motivated by proposed evolutionary scenarios in which black hole progenitor stars are rapidly rotating. For this reason, we do not include the ST dynamo (Spruit, 2002), which is commonly invoked in stellar models to enforce near-rigid rotation through efficient angular momentum coupling between stellar layers (Fuller and Ma, 2019; Fuller and Lu, 2022). The physical validity and efficiency of this mechanism remain uncertain, and several studies have questioned whether the magnetic instabilities required to sustain the ST dynamo can operate under realistic stellar conditions (e.g. Denissenkov and Pinsonneault, 2007; Zahn et al., 2007).
At the same time, it is widely recognized that some form of efficient angular momentum coupling must operate in stars in order to reproduce observational constraints, particularly those inferred from low- and intermediate-mass stellar remnants, such as the slow rotation of white dwarfs and neutron stars and the modest core–envelope differential rotation observed in evolved stars. Whether similarly efficient coupling operates in the very massive stars considered here, especially in the pair-instability regime, remains uncertain.
Models that adopt efficient angular momentum transport via the ST dynamo predict nearly non-rotating black holes (Marchant and Moriya, 2020), which are difficult to reconcile with the high component spins inferred for GW231123, potentially pointing to weaker internal angular momentum transport in its progenitor. [^1] We note that alternative interpretations suggest the nominal spin inference itself may be biased (Ray et al., 2025). We therefore allow for differential rotation and spin retention in our models, which both enables high-spin remnants and directly impacts the onset of pair-instability by shifting the upper edge of the mass gap to higher masses. Demonstrating that a rapidly spinning primary can still form above the mass gap strengthens the plausibility of a stellar-origin scenario involving photodisintegration collapse.
@@
Figure 1: Final BH mass as a function of initial rotation with initial helium core mass $160\rm{\rm M}_{\odot}$, for several values of the $^{12}\mathrm{C}(\alpha,\gamma)^{16}\mathrm{O}$ rate. The top panel shows results as a function of the initial stellar rotation $\Omega$ in units of the critical value $\Omega_{\rm crit}$; the bottom panel shows results as a function of the BH spin $\chi$. Stars with faster rotation and smaller $^{12}\mathrm{C}(\alpha,\gamma)^{16}\mathrm{O}$ reaction rate undergo PISN and are not shown in the figure. The shaded contours indicate the $90\%$ credible intervals of the primary mass (orange) and spin (purple) in GW231123 (Abac and others, 2025)
@@
Some simulations feature spin parameters exceeding unity at core-collapse. Although the BH may form with maximal spin, further accretion can reduce its spin if the added mass is not accompanied by sufficient angular momentum, or if excess angular momentum is expelled by magnetohydrodynamic feedback (e.g. McKinney et al., 2012)). To estimate the final spin of the BH, we adopt a version of the model developed by Batta and Ramirez-Ruiz (2019), consistent with previous work (Marchant and Moriya, 2020). We assume that the innermost $3{\rm M}_{\odot}$ of the progenitor collapses promptly into a BH with maximal spin. We have explicitly verified that the results are insensitive to the size of this core, see appendix A for more details. The remainder of the star is assumed to either fall directly into the BH or accrete through a disk, delivering additional mass and angular momentum. Accretion is regulated by feedback, ensuring that the final spin does not exceed the Kerr limit ($\chi\leq 1$).
4 Results
4.1 Impact of rotation on the black hole mass gap